OpenStudy (anonymous):

Shana uses a probability simulator to pull colored stars from a bag 60 times and to flip a coin 60 times. The results are shown in the tables below: Color of Star Number of Times Pulled Pink 17 Yellow 19 Violet 24 Heads Tails 26 34 Using Shana's simulation, what is the probability of pulling a pink star and the coin landing on tails?

Mehek (mehek14):

how many times was pink pulled out?

OpenStudy (anonymous):

A. fraction 578 over 3,600 B. fraction 646 over 3,600 C. fraction 34 over 60 D. fraction 51 over 60

OpenStudy (anonymous):

17

OpenStudy (anonymous):

@Mehek14 17

Mehek (mehek14):

how many total stars?

OpenStudy (anonymous):

dont know it says that they drew them out 60 times :/

Mehek (mehek14):

correct so \(\dfrac{17}{60}\) pink stars

Mehek (mehek14):

now \(\dfrac{tails}{total~fips}\)

OpenStudy (anonymous):

\[\frac{ 34 }{ 60}\]

Mehek (mehek14):

now multiply \(\dfrac{17}{60}*\dfrac{34}{60}=answer\)

OpenStudy (anonymous):

\[\frac{ 578 }{ 3600 }\]

OpenStudy (anonymous):

?

Mehek (mehek14):

so A \(\large{\ddot\smile}\)